OpenStudy (anikate):

find the exact value of cos(135) and sin(135)

OpenStudy (anikate):

@Calcmathlete

OpenStudy (anikate):

@ahaines14 @seattle12345 @sami-21 @Callisto

OpenStudy (anikate):

@Calcmathlete

OpenStudy (anikate):

help!

OpenStudy (anikate):

calling all units we need help on math

OpenStudy (anikate):

hellp me

OpenStudy (anonymous):

lol...it's called the unit circle.

OpenStudy (anikate):

wait

OpenStudy (anonymous):

Do you remember the values for it?

OpenStudy (anikate):

how do we find out without using the unit circle?

OpenStudy (anonymous):

Your back!

OpenStudy (anikate):

what's on my back?

OpenStudy (anonymous):

Not really any other way at this point. A bit of advice, you probably should memorize the unit circle, or at least Quadrant I.

OpenStudy (anikate):

can you plz post it

OpenStudy (anonymous):

Got it?

OpenStudy (anikate):

how do I figure it out with the unit circle?

OpenStudy (anonymous):

cosine is the x-value and sine is the y-value.

OpenStudy (anonymous):

Its 2 root 2, isnt it?

OpenStudy (anonymous):

With a hypotenuse of 1

OpenStudy (anikate):

what does that mean calc?

OpenStudy (anonymous):

\[\cos 135º = -\frac{\sqrt{2}}{2}\]\[\sin 135º = \frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

Basically, the coordinates for 135º on the unit circle is \[(-\frac{\sqrt{2}}{2}, \frac{\sqrt{2}}{2})\]cosine is the x-value and sin is the y-value.

OpenStudy (anonymous):

right, because it is in the second quadrant, and you have to go back.707, then up .707

OpenStudy (anonymous):

Thats how I think of it

OpenStudy (anonymous):

Yeah. Also, @Topkart33 usually when you use the unit circle, you don't use a calculator because that defeats the purpose of using it..

OpenStudy (anikate):

will it always be cos and sin only what about tan?

OpenStudy (anonymous):

Remember that \[\tan = \frac{\sin}{\cos}\]

OpenStudy (anonymous):

Right, it just gives me a perspective of what the traingel looks like in relationship to where it is on the circle, I know it is root 2/2

OpenStudy (anonymous):

triangle*

OpenStudy (anikate):

is there a numonis device to remember all the formulas?

OpenStudy (anonymous):

For the sin, cos, and trig part?

OpenStudy (anonymous):

*tan

OpenStudy (anikate):

no for sn=a1(n^r-1) etc.

OpenStudy (anikate):

calcmath what time is it in your place?

OpenStudy (anonymous):

ok lets do it \[\Large \cos(90+45)=-\sin(45)=-\frac{1}{\sqrt{2}}\]

OpenStudy (anonymous):

Well, I don't have one for the sequences and series ones, but I never really needed it :/ It is 10:28

OpenStudy (anikate):

oh same here

OpenStudy (anonymous):

now for sin \[\Large \sin(90+45)=\cos(45)=\frac{1}{\sqrt{2}}\]

OpenStudy (anikate):

cos is x and is y, right?

OpenStudy (anonymous):

@Anikate did you get it ?

OpenStudy (anikate):

and sin is y right?

OpenStudy (anonymous):

10:28 PM?

OpenStudy (anonymous):

Now 10:31pm, yeah that is the proof of it

OpenStudy (anonymous):

@Anikate ??

OpenStudy (anikate):

cos is x and sin is y right?

OpenStudy (anikate):

someone answer me

OpenStudy (anonymous):

Yes. cos is x and sin is y.

OpenStudy (anonymous):

@Anikate what is the Question can you elaborate ? i did not get your question you want to find the values of sin(135) and cos(135) correct?

OpenStudy (anonymous):

Do you understand? You can also work it out if you want.